what is the factional change in the resistance of an iron filament when its temperature changes from 25 to 50 degrees celcius?
1
Expert's answer
2020-04-06T09:14:13-0400

The temperature dependence of resistance is given by equation

"R(t_2)=R(t_1)(1+\alpha (t_2-t_1))."

The fractional change in the resistance

"\frac{R(t_2)-R(t_1)}{R(t_1)}=\alpha (t_2-t_1)\\
=0.00651\times (50-25)=0.16."